OpenStudy (anonymous):

Hey guys im studying for a final and i would really know if you could help me with this problem . An emission ray has a frequency of 5.10 × 10¹Ê Hz. Given that the speed of light is 2.998 × 10§ m/s, what is the wavelength of the ray?

OpenStudy (unklerhaukus):

\[c=f\lambda\qquad\qquad\rightarrow\qquad\lambda=\frac cf\] \[f=5.10\times10^{10} \text {Hz}\qquad\qquad c=2.998\times10^8 \frac{\text m}{\text s}\] \[\lambda=\frac{2.998\times10^8 \frac{\text m}{\text s}}{5.10\times10^{10} \text {Hz}}=\cdots\frac{[\text m/\text s]}{[1/\text s]}=\cdots [\text m]\]

OpenStudy (unklerhaukus):

for some reason i guessed that Ê=0

OpenStudy (anonymous):

thanks love :) i really have trouble with my chemistry class :( i try so hard but i never get them v__V but thanks :) is there any way you can explain to me how you really understood the problem.

OpenStudy (unklerhaukus):

i was picturing the photon

OpenStudy (anonymous):

@UnkleRhaukus so what was the answer?